A pair of shoes costs £5 more than twice the cost of a tie. If Grant buys two pairs of shoes and three ties, the total cost is £255. Find the cost of:
a) a tie
b) a pair of shoes

Answer by harpazo(655)

One pair of shoes = 2x + 5

One tie = x

2(2x + 5) + 3x = 255

4x + 10 + 3x = 255

7x + 10 = 255

7x = 255- 10

7x = 245

x = 245/7


x = 35

One tie sells for 35£.

One pair of shoes sells for

2(35) + 5 = 75£.

I hope this helps.